OpenStudy (anonymous):

Prove that E(X-µ)2 = E(X2) – µ2 Prove that V (aX)=a^2 V(X)

OpenStudy (anonymous):

what is E and V?

OpenStudy (anonymous):

hey girl

OpenStudy (anonymous):

i believe expectancy and variance

OpenStudy (anonymous):

oh, sry can't help in this :(

OpenStudy (zarkon):

let \(E(X)=\mu\) then \(E[(X-\mu)^2]=E[X^2-2\mu X+\mu^2]\) \(=E[X^2]-2\mu E[X]+E[\mu^2]=E[X^2]-2\mu\mu-\mu^2\) \(=E[X^2]-\mu^2\)

OpenStudy (zarkon):

use \(V[X]=E[(X-\mu)^2]\) and the first part to get the second part